Electric charge is distributed over the disk
x^2+y^2≤2 so that the charge density at (x,y) is σ(x,y)=17+x^2+y^2 coulombs per square meter.
Find the total charge on the disk.
1
Expert's answer
2020-11-09T19:58:01-0500

The distribution of electric charge can be described in polar coordinates as:

"x=rcos\theta, y=rsin\theta, r^2= x^2+y^2" hence

"0\leq r\leq \sqrt{2}" and "0 \leq \theta \leq 2\pi"


The total charge on the disk would the the double integral of the density function:

"\int_0^{2\pi}\int_0^{\sqrt{2}}(r^2+17)rdrd\theta=\int_0^{2\pi}\int_0^{\sqrt{2}}(r^3+17r)drd\theta="

"=\int_0^{2\pi}(\frac{r^4}{4}+\frac{17r^2}{2})\mid_0^{\sqrt{2}}d\theta=\int_0^{2\pi}(\frac{4}{4}+\frac{17\cdot2}{2}-\frac{0}{4}-\frac{17\cdot0}{2})d\theta="

"=\int_0^{2\pi}18d\theta=18\theta\mid_0^{2\pi}=18\cdot2\pi-18\cdot0=36\pi"


Answer: "36\pi"